Suppose that investment expenditures increases by $300 billion in a closed and private economy (no government or foreign trade). Assume further that households have a marginal propensity to consume of 75 percent. What will be the final cumulative impact on spending?

$900 billion

$225 billion

$375 billion

$525 billion
Answer #1

Investment expenditure increases by $300 billion in a closed and private economy and here no government and foreign trade. Household has marginal propensity to consume 75% i.e 0.75. The final and cumulative impact depend on spending multiplier. The spending multiplier is 1/1-MPC. The cumulative impact (1/1-MPC)* change in investment expenditure. Therefore 1/1-MPC = 1/1-0.75 = 1/0.25 = 4. The cumulative impact will be equal to 4*change in investment expenditure = 4*$300 = $1200.

None of the option is true. The final and cumulative impact will be $1200.